Status Report

NASA Space Station On-Orbit Status 24 Nov 2003

By SpaceRef Editor
November 25, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 24 Nov 2003
iss

All ISS systems continue to function nominally, except as noted previously or below. Underway: Week 5 of Expedition 8.

After station inspection and morning hygiene, before breakfast and first exercise, CDR/SO Michael Foale and FE Alexander Kaleri performed their second session of the periodic Russian biomedical assessments PZEh-MO-8 (body mass measurement) and PZEh-MO-7 (calf volume measurement). [Kaleri set up the BMM mass measuring device, which uses calibrated springs to determine the subject’s mass in weightless space, and stowed it away after the tests. Calf measurements (left leg only) are taken with the ISOG device, a custom-sewn fabric cuff that fits over the calf, using the knee and lower foot as fixed reference points, to provide a rough index of deconditioning in zero-G and effectiveness of countermeasures.]

The remainder of today’s schedule was largely replanned around the high-priority data collection of contingency TVIS treadmill operating modes (ref.: yesterday’s On-Orbit Status report).

The crew first removed an IWIS (interim wireless instrumentation system) sensor package from a location in the Lab and reinstalled it near the TVIS in the Service Module. Kaleri also had to recharge the batteries of the Russian DVCAM PD-1P and -150P camcorders for the subsequent videotaping. [The IWIS components consist of an accelerometer for measuring structural vibrations and its wireless (radio-equipped) remote sensing unit (RSU). The IWIS setup was required to obtain loads imposed on the station structure when the treadmill was operated in several optional contingency modes without use of the roll-stabilizing gyroscope.]

The testing itself plus accompanying data gathering began at about 10:00am EST, performed by both crewmembers for several hours. [The uplinked procedure called for each crewmember to perform three minutes walking plus three running for getting stabilized with or without handrails and SPDs (subject positioning devices), then collect approximately two minutes of data in each of 4-5 alternate session modes. Along with the dynamic loads measurements, the crew took video recordings of the walking and running tests on the treadmill on four tapes, two from stationary cameras and two from handheld cameras.]

Later, the IWIS equipment was removed again, and the vibration data collected from SM and Node were downloaded to the file server for analysis by ground specialists, along with the camcorder video. 

Alex Kaleri worked in the FGB, removing two brackets behind a panel for clearing an IMV (intermodular ventilation) air duct.
 
Good news from Moscow:  The Russian VELO cycle ergometer is back in operation, and Kaleri already performed an exercise on it today. Along with the U.S. CEVIS, this provides the crew now with two bikes for aerobic exercise. [The restoration did not require the planned disassembly of the control module, but this task is retained in the plan for the time being.]

Imagery of the recent stowage equipment relocation from the Node to PMA-2, PMA-1 and behind two of five available zero-G storage racks (ZSRs) indicates that all five fire ports in the Node are now fully accessible. [Access to 19 other U.S. segment (USOS) fire ports has not been an issue.]

Sasha completed the daily routine maintenance of the SM’s SOZh life support system, including ASU toilet facilities, and the preparation of the “delta” file for updating the IMS today. Mike conducted the regular routine status checkup of running ISS-8 payloads.

Before sleep time tonight, the FE will set up the test equipment for the periodic Russian MO-10 “Hematokrit” testing scheduled for tomorrow. [MO-10 measures the hematocrit (red cell mass) value of the blood (as a well-known phenomenon of space flight, red blood cell mass {normal range: 30-45%} tends to go down over time).]

Due to time constraints, the planned preparation of the required equipment for the upcoming recharging of the Motorola-9505 Iridium satphone in the Soyuz TMA-3 descent module was deferred from today to tomorrow. [Prior to charging, the phone is to be placed in a single CTB (crew transfer bag), which then is packed into a triple CTB (prepared today as per MCC-H instructions). A checkup of the recharge process is to be conducted every 30 minutes without removing the phone from the CTBs. After recharging is completed, power will be disconnected and the phone removed from the CTBs and restowed in the Iridium SSSP kit without removing the battery from the phone. The kit will then be placed in the crew procedures container of the Soyuz descent module.]

The date for the attitude transition maneuver from the current sun-pointing XPOP (x-axis perpendicular to orbit plane) to earth-oriented LVLH (local vertical/local horizontal) has been moved forward from the Thanksgiving weekend (11/28) to next Wednesday (11/26), with concurrence from Moscow.

The NASA/RSA meeting to discuss plans for the two-person Orlan-spacewalk early next year is now scheduled for 12/1 at JSC, with Russian specialists in attendance.

Moscow reports that a relocation of the Soyuz TMA-3 from the DC-1 docking module to the FGB nadir port is not planned in the current Increment. [MCC-H has put together a list of preferable and avoidable dates in case the relocation should become necessary after all.]

Tomorrow, MCC-H starts a two-week test of transmitting video (received from the station via Ku-band) from White Sands to JSC and MSFC via landline instead of via domestic comsat. [White Sands to JSC is already using landline. The test will extend this to Huntsville. Landline is more economic.]

Today’s CEO (crew earth observation) targets, limited in the current XPOP attitude by flight rule constraints on the use of the Lab nadir/science window, were Inland deltas, S Sudan (Dynamic event. Weather remains clear in this remote, war-torn region where geological research stopped decades ago. Sunglint images left of track were requested. A series of non-optimum handheld images [cloudy and distant] has been used to reconstruct the landscape of this previously unstudied basin, west of the Nile and Sudd swamps. This landscape is likely a new analog for known oil-rich rocks buried far beneath the surface. Better images are requested), Andean foothills fires (Dynamic event. Terra satellite images show hot spots mainly along the eastern Andean foothills, with some fires on the wide plains west of the Andes. Looking left and right for smoke plumes), West African fires (looking right of track along the axis of the Sahel savanna zone, in which numerous fires are reported), Lower Amazon River Basin (looking mainly left for estuary. Dry weather in the past weeks has resulted in fires, which are reported in this equatorial forest), and Lima, Peru (nadir pass. Looking inland a few miles for the city center).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:40pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 19A. Vozdukh CO2 scrubber is On (in Manual Mode 5/3). U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. MCA (major constituents analyzer) is off (in Life Extending Mode). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 27.5; ppO2 (mmHg) — 152.9; ppCO2 (mmHg) — 4.0.
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 21.1.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 23.0.
  • Node:  Pressure (mmHg) — 743.89; temperature (deg C) — 24.4 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 745.42; temperature (deg C) — 24.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 745.62; temperature (deg C) — 28.0; shell heater temp (deg C) — 24.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 25.5
  • PMA-2:  Shell heater temp (deg C) — 14.2

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in biased Autotrack mode (suntracking). 
  • SM batteries:  Battery #8, formerly known as #7, is still in slot #8 for troubleshooting; all other batteries (7) are in “Partial Charge” mode (batteries #1 and #3 are degraded).  
  • FGB batteries:  Batteries #1 is off; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-2 MDM is prime, C&C-1 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational (all lanes reintegrated 11/5).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational; string #3 dropped out 10/22.

Propulsion System:

  • Total propellant load available: 3704 kg (8166 lb) as of 11/13  [SM(755) + FGB(2597) + Progress M(352) + Progress M-1(0)]. (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime, IAC-1 is suspect).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:27am EST [= epoch]):

  • Mean altitude — 374.4 km
  • Apogee — 378.3 km
  • Perigee — 370.5 km
  • Period — 92.0 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0005819
  • Orbits per 24-hr. day — 15.64
  • Mean altitude loss in last 24 hours — 120 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 28606
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html

SpaceRef staff editor.